OpenStudy (anonymous):

Solve for x in 2x^2 = 8

OpenStudy (anonymous):

First divide both sides by 2 x^2 = 4 and what is next?

OpenStudy (anonymous):

2x^2 = 8 x^2 = 4 x=2

OpenStudy (anonymous):

+-2

OpenStudy (anonymous):

almost, we have two answers ..

OpenStudy (anonymous):

x = +2 and -2